# Robust Market Interventions

### v1

@@ 2.1.2. Supply side
There is a firm associated with each product: Firm $i$ produces good $i$. Firms play a simultaneous pricing game; each firm chooses $\tilde{p}_{i}\geq 0$. For any realized profile of prices $\tilde{\bm{p}}$, firm $i$’s profit is
@@
From now on, we confine attention to cost perturbations within the set discussed in 1, and when we refer to an equilibrium at any cost profile, we mean the locally unique one entailed by the above assumption.
2.2. Interventions and outcomes
An authority—an institution overseeing a marketplace—can intervene on the market. For concreteness, we focus on a canonical set of interventions: per-unit subsidies and taxes. For a consumption profile $\tilde{\bm{q}}$, a per-unit subsidy intervention
@@ 2.2. Interventions and outcomes
consists of a transfer $\sigma_{i}\tilde{q}_{i}$ from the authority to firm $i$; a positive $\sigma_{i}$ corresponds to a subsidy to firm $i$, while a negative $\sigma_{i}$ corresponds to a tax. The authority’s net spending associated with such an intervention is denoted by $S=\bm{\sigma}\cdot\tilde{\bm{q}}$. Recalling 1, the set of feasible interventions is [^9] We restrict attention to deterministic interventions but this is immaterial to our results.
@@
We evaluate the effect of an intervention $\bm{\sigma}$ on an outcome variable $Y$ by its first derivative. Formally, the first-order effect on any given outcome variable $Y$ of changing subsidies in the direction $\bm{\sigma}$ is defined by [^10] We often omit the subscript $\bm{\sigma}$ when there is no ambiguity about the relevant intervention.
@@
2.3. A locally linear oligopoly: Market states and signals
The authority receives a noisy signal about the structure of the market and, using this information, wants to implement interventions that robustly increase standard measures of surplus.
Before giving a general definition of this problem and of robust interventions, we introduce a special case—one with demand functions that are linear in a neighborhood of the status quo equilibrium, as captured by the following assumption.
Assumption 2 (Local linearity of demand).
This setting is highly tractable and, as we will see, allows us to formulate an interesting robust intervention problem. We discuss how to extend the analysis for the case of non-linear demand in Section 7.4. We now describe what the authority would ideally like to know, the signals it actually has, and then define the notion of a robust intervention.
${q}_{i}(\bm{p})=-\frac{\partial q_{i}}{\partial p_{i}}(\bm{p})({p}_{i}-{c}_{i}).$
Local linearity implies that $\frac{\partial q_{i}}{\partial p_{i}}(\bm{p})$ is constant when prices change locally around $\bm{p}^{0}$. Hence, we can replace the partial derivative in the above equation with $\frac{\partial q_{i}}{\partial p_{i}}(\bm{p}^{0})$. By strict concavity of the consumer utility functions, this is a negative number; from now on, we maintain a normalization (by choosing suitable units in which to express the quantity produced by firm $i$) that $\frac{\partial q_{i}}{\partial p_{i}}(\bm{p}^{0})=-1$ (see Section A.1). After this normalization, equilibrium behavior is summarized by the following system of equations:
$\bm{q}(\bm{p})=\bm{p}-\bm{c}.$ (4)
$[\bm{I}-\bm{D}]\dot{\bm{p}}=-\bm{\sigma},$ (5)
where $\bm{\sigma}=\bm{c}-\bm{c}^{0}$ is the intervention (i.e., tax or subsidy offered by the authority), $\dot{\bm{p}}$ is the derivative of $\bm{p}$ in the direction of $\bm{\sigma}$ (see eq. 3), and $\bm{D}=\bm{D}(\bm{p}^{0})$ is the Slutsky matrix (in the normalized units):
$D_{ij}={\frac{\partial q_{i}}{\partial p_{j}}}(\bm{p}^{0}).$
For $i\neq j$, if $D_{ij}>0$ (resp. $D_{ij}<0$) then, around the equilibrium, products $i$ and $j$ are substitutes (resp. complements). [^11] In general, the matrix of demand derivatives need not be the same as the Slutsky matrix, but under quasilinearity the wealth effect is zero and the two coincide (Nocke and Schutz, 2017). Note also that quantity changes following the intervention $\bm{\sigma}$ are pinned down by
$\dot{\bm{q}}=\bm{D}\dot{\bm{p}}.$ (6)
The local linearity of demand implies that comparative statics of prices and quantities are fully determined by the Slutsky matrix $\bm{D}$. This implies the following fact, which we will establish and discuss in our analysis below.

### v2

@@ 2.1.2. Supply side
There is a firm associated with each product: Firm $i$ produces product $i$. Firms play a simultaneous pricing game; each firm chooses $\tilde{p}_{i}\geq 0$. For any realized profile of prices $\tilde{\bm{p}}$, firm $i$’s profit is
@@
From now on, we confine attention to cost perturbations within the set discussed in 1, and when we refer to an equilibrium at any cost profile, we mean the locally unique one entailed by this assumption.
2.2. Interventions and their effects
An authority—an institution that oversees a marketplace—can intervene in the market. We focus on a canonical set of interventions: per-unit subsidies and taxes. For a consumption profile $\tilde{\bm{q}}$, a per-unit subsidy intervention
@@ 2.2. Interventions and their effects
consists of a transfer $\sigma_{i}\tilde{q}_{i}$ from the authority to firm $i$; a positive $\sigma_{i}$ corresponds to a subsidy to firm $i$, while a negative $\sigma_{i}$ corresponds to a tax.
We will focus on demand functions that are linear in a neighborhood of the status quo equilibrium, as captured by the following assumption.
Assumption 2 (Linearity of demand locally).
This assumption facilitates our analysis of interventions, yielding simple formulas for comparative statics. It also captures much of the economics of our robust interventions problem, despite its simplicity. We discuss how to extend the analysis for the case of non-linear demand in Section 7.4.
${q}_{i}(\bm{p})=-\frac{\partial q_{i}}{\partial p_{i}}(\bm{p})({p}_{i}-{c}_{i}).$
The linearity assumption implies that $\frac{\partial q_{i}}{\partial p_{i}}(\bm{p})$ remains constant when prices change locally around $\bm{p}^{0}$. Hence, we can replace the $\bm{p}$-dependent partial derivative in the above equation with the constant $\frac{\partial q_{i}}{\partial p_{i}}(\bm{p}^{0})$. By strict concavity of the consumer utility functions, this is a negative number; from now on, we maintain a normalization (by choosing suitable units in which to express the quantity produced by firm $i$) that $\frac{\partial q_{i}}{\partial p_{i}}(\bm{p}^{0})=-1$ (see Section A.1). After this normalization, equilibrium behavior is summarized by the following system of equations:
$\bm{q}(\bm{p})=\bm{p}-\bm{c}.$ (2)
$[\bm{I}-\bm{D}]\dot{\bm{p}}=-\bm{\sigma},$ (3)
where $\bm{\sigma}=\bm{c}^{0}-\bm{c}$ is the intervention (i.e., tax or subsidy offered by the authority), $\dot{\bm{p}}$ is the derivative of $\bm{p}$ in the direction of $\bm{\sigma}$ (see eq. 6), and $\bm{D}=\bm{D}(\bm{p}^{0})$ is the Slutsky matrix (in the normalized units):
$D_{ij}={\frac{\partial q_{i}}{\partial p_{j}}}(\bm{p}^{0}).$
For $i\neq j$, if $D_{ij}>0$ (resp. $D_{ij}<0$) then, around the equilibrium, products $i$ and $j$ are substitutes (resp. complements). [^10] In general, the matrix of derivatives of Marshallian demand need not be the same as the Slutsky matrix (which works with compensated demand). However, in this demand system, the wealth effect is zero due to the fact that the goods utility and money are additively separable. Thus, the two matrices coincide (Nocke and Schutz, 2017), and so we use the term “Slutsky matrix” throughout. Note also that quantity changes following the intervention $\bm{\sigma}$ are pinned down by
$\dot{\bm{q}}=\bm{D}\dot{\bm{p}}.$ (4)
The local linearity of demand implies that comparative statics of prices and quantities are fully determined by the Slutsky matrix $\bm{D}$. We note that $\bm{D}$ satisfies the following property (Nocke and Schutz, 2017).
Property NSD.
Finally, the following helpful normalization is without loss of generality—it holds by suitably adjusting the units of the numeraire $m$.
Assumption 3.
The quantity vector’s Euclidean norm $\|\bm{q}^{0}\|$ is at most 1.
In what follows, we maintain Assumptions 1–3 unless stated otherwise.
2.3. Surpluses
The authority’s net spending associated with an intervention $\bm{\sigma}$ is denoted by $S=\bm{\sigma}\cdot\tilde{\bm{q}}$. Recalling 1, we define the set of feasible interventions as [^11] We restrict attention to deterministic interventions but this is immaterial to our results.
@@ 2.3. Surpluses
We evaluate the effect of an intervention $\bm{\sigma}$ on an outcome variable $Y$ by its first derivative. Formally, the first-order effect on any outcome variable $Y$ (e.g., $Y=C$ or $P$) of changing subsidies in the direction $\bm{\sigma}$ is defined by [^12] We often omit the subscript $\bm{\sigma}$ when there is no ambiguity about the relevant intervention.
@@
Fixing the parameters of the economy, the set of possible surplus outcomes is defined to be the set of tuples
$\{(\dot{C}_{\bm{\sigma}},\dot{P}_{\bm{\sigma}},\dot{S}_{\bm{\sigma}}):\bm{\sigma}\in\mathbb{R}^{n}\}$
of surplus outcomes corresponding to some intervention. The following proposition characterizes the possible surplus outcomes.
Proposition 1.
For generic $\bm{q}^{0}$ and any $\bm{D}$, the following hold:
  1. (1)
    A surplus outcome $(\dot{C},\dot{P},\dot{S})$ is possible if and only if it satisfies
    $\dot{C}+\frac{1}{2}\dot{P}=\dot{S}.$ (7)
  2. (2)
    If $(\dot{C},\dot{P},\dot{S})$ is possible and $\dot{C}\geq 0$, then
    $\dot{C}+\dot{P}\leq 2\dot{S}.$
Proof sketch..
Consumer surplus satisfies the Marshallian formula $\dot{C}=-\bm{q}^{0}\cdot\dot{\bm{p}}$. Meanwhile, $\dot{P}=\bm{q}^{0}\cdot(\dot{\bm{p}}+\bm{\sigma})+(\bm{p}-\bm{c})\cdot\dot{\bm{q}}=2\bm{q}^{0}\cdot\dot{\bm{q}}$, where we have used the equilibrium condition $\bm{p}-\bm{c}={\bm{q}}$ from (2). Finally, $\dot{S}=-{\bm{q}}^{0}\cdot\dot{\bm{c}}={\bm{q}}^{0}\cdot\dot{\bm{\sigma}}$ by definition. This establishes that any $(\dot{C}_{\bm{\sigma}},\dot{P}_{\bm{\sigma}},\dot{S}_{\bm{\sigma}})$ satisfies (7). The proof that every tuple satisfying () arises from some $\bm{\sigma}$ is more involved and is found in Appendix B. The second part follows from the first. ∎
Part (1) says that, for a given level of spending $\dot{S}$, the set of possible outcomes is a line in $(\dot{P},\dot{C})$ space. This “Pareto frontier” tells us what is possible in principle, and in particular implies that market surplus cannot be increased by more than twice the level of expenditure without reducing consumer surplus.
While this result characterizes what post-intervention outcomes are possible in equilibrium, it does not discuss how to attain them. The following lemma is an important input in the answer to this question, and is used in the proof of the “if” direction of 1(1).
Lemma 1.
For any intervention $\bm{\sigma}$, spending is given by $\dot{S}_{\bm{\sigma}}=\bm{\sigma}\cdot\bm{q}^{0}$, while
$\dot{W}_{\bm{\sigma}}=\frac{1}{2}\dot{P}_{\bm{\sigma}}=(\bm{q}^{0})^{\mathsf{T}}\bm{D}[\bm{I}-\bm{D}]^{-1}\bm{\sigma}$ (8)
The proof works by combining the formulas in the proof of the above proposition with the formulas in (3) and (4) to make price and quantity derivatives explicit.
Formula (8) can be interpreted as a pass-through equation: entry $i$ of the row vector $\bm{w}^{\mathsf{T}}=(\bm{q}^{0})^{\mathsf{T}}\bm{D}[\bm{I}-\bm{D}]^{-1}$ gives the impact on total surplus of increasing the subsidy $\sigma_{i}$. The authority aims to achieve a desired total surplus effect $\dot{W}_{\bm{\sigma}}$, possibly subject to additional requirements, such as holding spending $\dot{S}_{\bm{\sigma}}$ constant.
2.4. The challenge
Matrix inverses such as $[\bm{I}-\bm{D}]^{-1}$ can be extremely sensitive to entries of $\bm{D}$. Therefore, without precise knowledge of $\bm{D}$ and $\bm{q}^{0}$, the authority may not be able to implement a desired point on the line defined by (7). For example, the authority may not be sure that a given intervention will increase total surplus ($\dot{W}_{\bm{\sigma}}>0$) rather than decrease it ($\dot{W}_{\bm{\sigma}}<0$). Indeed, it seems hard to justify the detailed study of comparative statics such as eq. 8 when $n$ is large without confronting the uncertainty about the ingredients of the formula an analyst or authority is likely to face.
These observations motivate the central question of this paper: Which interventions have surplus effects that can be predicted with confidence by an authority facing substantial uncertainty about market primitives?
2.5. The statistical framework
To formalize this question, we introduce a simple model of noisy observation of $\bm{D}$ and $\bm{q}^{0}$.
In the linear oligopoly environment, 1 establishes that to determine the first-order surplus effects of any intervention $\bm{\sigma}$, the only additional data needed is the tuple $(\bm{D},\bm{q}^{0})$, where $\bm{D}$ is a negative semidefinite matrix with diagonal entries $-1$ and $\bm{q}^{0}$ is a vector of norm at most $1$. We call such a tuple a market state and denote it by $\bm{\theta}$. The set of possible market states is denoted by $\bm{\Theta}$.
The authority receives a signal, denoted by $\widehat{\bm{\theta}}\in\widehat{\bm{\Theta}}$, about the market state. [^13] Section 7.1 discusses some practical examples. This signal consists of random variables
[…]
We will later detail assumptions on the error terms. For now, a canonical setting to keep in mind is one where all error draws are mean-zero and independent, with each error having a magnitude comparable to the underlying entry $D_{ij}$ or $q^{0}_{i}$. Note that the signal need not lie in the same set as the state; for example, under our assumptions, $\bm{D}$ is negative semidefinite, but the signal $\widehat{\bm{D}}$ might not be.
Let $\varphi_{\bm{\theta}}\in\Delta(\widehat{\bm{\Theta}})$ denote the probability measure over signals when the state is $\bm{\theta}$.
2.6. Robust intervention rules
The authority designs an intervention rule [^14] This should be measurable in a suitable sense, which is clear in our application.
$\bm{R}:\mathcal{T}\to\bm{\Sigma},$
prescribing an intervention $\bm{\sigma}\in\bm{\Sigma}$ for every possible signal $\widehat{\bm{\theta}}$. We now define a notion of such a rule robustly achieving a desired property.
A market outcome is a tuple $(\bm{\theta},\bm{\sigma})$ consisting of a market state and an intervention. We call this pair the outcome because it determines production, consumption, and transfers. A property is a measurable subset $\mathscr{P}\subseteq\bm{\Theta}\times\bm{\Sigma}$ of all possible outcomes. An important example of a property is increasing total surplus, given by (8):
$\mathscr{P}$ =$=\{(\bm{\theta},\bm{\sigma}):\dot{W}_{\bm{\sigma}}>0\}$
=$=\{((\bm{D},\bm{q}^{0}),\bm{\sigma}):(\bm{q}^{0})^{\mathsf{T}}\bm{D}[\bm{I}-\bm{D}]^{-1}\bm{\sigma}>0\}$
We are interested in understanding which properties can be achieved with high probability in all market states that the authority considers possible:
Definition 1.
An intervention rule $\bm{R}$ achieves a property $\mathscr{P}$ $\epsilon$–robustly if the following holds: For every $\bm{\theta}\in\bm{\Theta}$, we have
$\varphi_{\bm{\theta}}\left(\{\widehat{\bm{\theta}}:(\bm{\theta},\bm{R}(\widehat{\bm{\theta}}))\in\mathscr{P})\}\right)\geq 1-\epsilon.$
The only randomness in the definition is in the signal draw—recall $\varphi_{\bm{\theta}}$ is the distribution of the signal given the true state $\bm{\theta}$.
We close with some remarks on our modeling choices.
Remark 1.
The main parameter in the definition of robustness is the set $\bm{\Theta}$ of possible market states. If the authority has any prior $\mu$ over $\bm{\Theta}$, then, given conditional error distributions $(\varphi_{\bm{\theta}})_{\bm{\theta}}$ and a property $\mathscr{P}$ of interest, the probability $\mathbb{P}_{\mu}(\mathscr{P})$ can be computed under that prior. Requiring an intervention rule to achieve a property $\epsilon$-robustly is equivalent to requiring $\mathbb{P}_{\mu}(\mathscr{P})\geq\epsilon$ for every possible prior $\mu$.
A reason to focus on robust intervention rules is that for primitives such as Slutsky matrices involving a large number of goods, it is not at all obvious how an authority should specify a prior. A set $\bm{\Theta}$ can, in contrast, encode an assumption on the environment (e.g., $\bm{D}$ being negative semidefinite) without having to take a stand on prior probabilities.
Remark 2.
Notice that the definition of robust interventions does not depend on any details of the linear oligopoly environment. The definition can be applied in any environment where $\bm{\theta}$ is some state, $\bm{\sigma}$ is the authority’s choice, the pair $(\bm{\theta},\bm{\sigma})$ fully determines all outcomes of interest to the authority, and there is a known distribution over signals given states.
Nevertheless, the linear oligopoly game and the focus on first-order effects provide a useful starting point: a canonical environment in which rich spillovers can be fully specified by a familiar finite-dimensional object ($\bm{D}$). Moreover, formulas using only simple matrix operations suffice to describe surplus effects, and these facilitate tractability of statistical questions. Extensions to nonlinear environments are discussed in Section 7.4.
3. Recoverable structure and its economic implications
The main result of this paper says that under conditions on the set of possible market states $\bm{\Theta}$ and conditions on the distribution of errors, there are intervention rules that improve surplus robustly. The content of the result lies in specifying the assumptions on $\bm{\Theta}$ and error distributions. While these conditions are somewhat involved to state in full generality, a useful preview can be presented in an important class of examples, related to the classic stochastic block model in network theory.
3.1. A stochastic block model of demand
There is a fixed, finite set of product types, $M=\{1,2,\ldots,m\}$. The interactions of products are determined by their types, and given by entries of an $m$-by-$m$ type-level matrix $\bar{\bm{D}}$ satisfying Property NSD. Quantities are also determined by types, according to a vector $\bar{\bm{q}}\in\mathbb{R}^{m}$. Let the type of good $i$ be $k(i)\in M$ [^15] We let $N(t)$ be the number of products of type $t$, which depends on the total number of products $n$, and assume $N(t)/n$ is a convergent sequence as $n\to\infty$ for each $t\in M$. Finally, fixing $\gamma(n)\in(0,1]$, let
$D_{ij}=\begin{cases}\gamma(n)\bar{D}_{k(i)k(j)}&\text{$i\neq j$}\\ -1&\text{$i=j$}\end{cases}\quad\text{ and }\quad q_{i}^{0}=\bar{q}_{k(i)}.$ (9)
An explicit example of a stochastic block model appears in Section 5. The errors $\varepsilon_{i}$ in observing $\bm{q}^{0}$ and the errors $E_{ij}$ in observing $\bm{D}$ are i.i.d. and mean zero, with bounded support, satisfying $\operatorname{Var}[E_{ij}]<\overline{\sigma}$ and $\operatorname{Var}[\|\bm{\varepsilon}\|]<\overline{\varsigma}\|\bm{q}^{0}\|$ for some fixed real numbers $\overline{\sigma}$ and $\overline{\varsigma}$. This entails that errors in demand signals and quantity signals are of the same order of magnitude as the underlying parameters. Within this model, we have:
Proposition 2.
Fix a generic $(\bar{\bm{D}},\bar{\bm{q}})$ and assume $\gamma(n)n^{1/2}\to\infty$. For large enough $n$ and any spending level $s$, there is an intervention rule that robustly achieves a surplus outcome arbitrarily close to $(\dot{P},\dot{C},\dot{S})=(2s,0,s)$.
Note that 1(2) implies that the intervention rules guaranteed by 2 achieve the largest possible increase in total surplus subject to not reducing consumer surplus. It is also worth noting that the partitioning of goods into types need not be known in advance to achieve the target surplus outcome.
In this setting, the assumption that $\gamma(n)n^{1/2}$ is large ensures that some of the entries of $\bar{\bm{D}}$ can be recovered despite noisy observation. Notice, however, that other entries of $\bar{\bm{D}}$ may not be recoverable precisely. One extreme example of this occurs if some types have only a single good, or more generally a number of goods uniformly bounded in $n$. The substance of the result is that the information that can be recovered about $\bar{\bm{D}}$ suffices to robustly achieve the indicated outcome.
The result raises several questions: First, how is recoverable information used to design effective interventions, and what determines the limits of this strategy? More fundamentally, how can it be extended beyond the specific structure of the stochastic block model? The assumption that there are arbitrarily large [^16] When there are $|M|$ types of products, then some type must contain at least $n/|M|$ products. “blocks” of products whose exact relationships to other goods (the entries $D_{ij}$) are identical within type is restrictive. Are there more flexible structures that permit robust interventions—e.g., in cases where no entry of $\bm{D}$ can be recovered precisely? A third question that this result raises is whether outcomes that allocate surplus differently can be robustly achieved. The remainder of the paper addresses the three issues we have raised.
The next subsection presents the assumption on $\bm{\Theta}$, called recoverable structure, that underlies our analysis of the general robust intervention problem. We then present the key method—a spectral price theory decomposition—that enables our use of this assumption, and illustrate the ideas throughout in relation to the stochastic block model special case.
3.2. Recoverable structure
Recoverable structure imposes conditions on the pattern of complements and substitutes (which we will call interactions) among products, as summarized by the Slutsky matrix $\bm{D}$. It requires that there is a strong latent pattern of product interactions and, simultaneously, this latent structure “has enough correlation” with the vector of market quantities.
We now define this notion formally. A vector in $\mathbb{R}^{n}$ describes a bundle of products. Given $\bm{D}$, we are interested in the subspace of bundles spanned by eigenvectors of $\bm{D}$ with large eigenvalues. Formally, let $\mathcal{L}(\bm{D},M)\subseteq\mathbb{R}^{n}$ be the subspace of the bundle space spanned by the eigenvectors of $\bm{D}$ with eigenvalues at least $M$ in absolute value.
Definition 2.
To understand this definition, note that a Slutsky matrix, by virtue of being symmetric, can be orthogonally diagonalized: it can be written as a linear combination of orthogonal rank-one matrices:
where $\lambda_{1},\lambda_{2},\ldots,\lambda_{n}$ are the eigenvalues of $\bm{D}$ (which are nonpositive numbers because $\bm{D}$ is negative semidefinite), ordered from greatest to least in absolute value, and $(\bm{u}^{1},\dots,\bm{u}^{n})$ is a corresponding basis of orthonormal eigenvectors. All the summands are orthogonal to each other, and $|\lambda_{\ell}|$ is the norm of the contribution of the corresponding summand.
Having market states with $(M,\delta)$-recoverable structure means that (i) $\bm{D}$ has eigenvectors with eigenvalues larger than $M$ in absolute value, ensuring the first summation is nonzero, and (ii) the vectors $\bm{u}^{\ell}$ in that sum can jointly account for a non-negligible portion of the status quo quantities. We will call the eigenvectors $\bm{u}^{\ell}$ associated with the eigenvalues such that $|\lambda_{\ell}|\geq M$ the top eigenvectors.
As we will detail later, if we set $M$ to be larger than the norm of the noise $\bm{E}$ in the signal of $\bm{D}$, then condition (i) ensures that the space of top eigenvectors of $\bm{D}$ can be estimated precisely. We now explain what recoverable structure with a large $M$ means economically before explaining the intuition behind requirement (ii), which we defer to Section 3.3.
3.2.1. Recoverable structure and large-scale complementarities
We start with a simple example.
Example 1.
Consider the special case where there is only one type of product, i.e., $M={1}$ and $\bar{D}_{k(i),k(j)}=k\in(0,1)$ and $\gamma(n)=1$. Then
$D_{ij}=\begin{cases}-1&\text{if }i=j,\\ -k&\text{if }i\neq j.\end{cases}.$
This can be written as
[…]
where $\bm{I}$ is the identity matrix and $\mathbf{1}$ is the all-ones vector. The eigenvalues of $k$ are $k-1$ with multiplicity $n-1$ and $-k(n-1)-1$ with multiplicity $1$. Thus, if $k>M/(n-1)$, and $\sum_{i}q^{0}_{i}>\delta$, the market state $(\bm{D},\bm{q}^{0})$ has $(M,\delta)$-recoverable structure.
Notice that negative off-diagonal entries in $\bm{D}$ correspond to product complementarities. Thus, even a small (entry-wise) amount of complementarity embedded in the demand system can generate a large eigenvalue and recoverable structure.
Recoverable structure is closely related to large-scale complementarities more generally. To see this, let $\|\bm{x}\|$ be the Euclidean norm of vector $\bm{x}$ and note that:
$|\lambda_{1}|=\sup_{\dot{\bm{p}}\neq\bm{0}}\frac{\|\bm{D}\dot{\bm{p}}\|}{\|\dot{\bm{p}}\|}=\sup_{\dot{\bm{p}}\neq\bm{0}}\frac{\|\dot{\bm{q}}\|}{\|\dot{\bm{p}}\|},$
where the first equality is the well-known Courant–Fischer characterization of the spectral radius of a symmetric matrix, and the second equality follows from of equilibrium condition, $\dot{\bm{q}}=-\bm{D}\dot{\bm{p}}$. The solution to the maximization problem over $\dot{\bm{p}}$ is to choose a change in price equal to the dominant eigenvector of $\bm{D}$, denoted by $\bm{u}^{1}$, which is the eigenvector associated with the largest eigenvalue in absolute value. Thus, the largest eigenvalue measures how much a price shock can be amplified in terms of its effect on demand and the corresponding eigenvector gives this extremal price shock achieving this effect.
This gives us a useful perspective on the meaning of a large $|\lambda_{1}|$. Note that if there were no demand spillovers across products (i.e., all products were independent), then by our normalization we would have $\bm{D}=-\bm{I}$ and each good’s quantity would change by an amount equal to the price change. When there are demand spillovers across products, and we change prices in the direction of $\bm{u}^{1}$, each good’s quantity changes by $-|\lambda_{1}|$ times the price change, a large negative multiple. Hence, the case $|\lambda_{1}|>1$ indicates that there are price changes where the downward effect on quantities is larger than in the independent case. This corresponds to complementarities, in that reductions in demand reinforce one another. [^17] In contrast, when $|\lambda_{1}|<1$, the effect on demand is lower than in the independent case, corresponding to an economy where products are globally substitutes.
This interpretation of the dominant eigenvector and its eigenvalue extends to the other eigenvectors and eigenvalues of $\bm{D}$. Let $\bm{u}^{\ell}$ be the eigenvector associated with the $\ell^{\text{th}}$ largest eigenvalue, denoted $\lambda_{\ell}$. A direct implication of the Courant–Fischer theorem is that the price change that maximizes the change in quantities relative to the change in price across all price changes orthogonal to ${\bm{u}^{1},\bm{u}^{2},\ldots,\bm{u}^{\ell-1}}$ is exactly $\bm{u}^{\ell}$, and
$\sup_{\dot{\bm{p}}\perp{\bm{u}^{x}}{x=1}^{\ell-1}}\frac{\|\dot{\bm{q}}\|}{\|\dot{\bm{p}}\|}=|\lambda_{\ell}|.$
Hence, the spectral decomposition of $\bm{D}$ captures a set of $n$ orthogonal price changes in the economy, each representing the maximum induced change in quantity that is feasible within the system given the orthogonality constraint.
3.2.2. Application to the stochastic block model
We now make explicit the relationship between the general recoverable structure condition and the stochastic block model. In the stochastic block model, fixing $\bar{\bm{D}}$, $\bar{\bm{q}}$, and a sequence $\gamma(n)$, the set of market states is the set of all $(\bm{D},\bm{q}^{0})$ (for any number of firms) satisfying the block specification (9) for some function $k$ partitioning products into types.